OpenStudy (anonymous):

14 less than one-half of a number is 2 more than the number

OpenStudy (amonoconnor):

(.5x) - 14 = x + 2

OpenStudy (anonymous):

im supposed to find the value of x

OpenStudy (amonoconnor):

Yeah... Want help?

OpenStudy (anonymous):

yeah

OpenStudy (amonoconnor):

K. .5x - 14 = x + 2 + 14 _______________ .5x = x + 16 -x ____________ (-.5x) = 16 ___________ (-.5) X = 16/(-.5) = -32

OpenStudy (anonymous):

alright , i get it

OpenStudy (amonoconnor):

:) cool, next?

OpenStudy (anonymous):

x+10=27 , solve equation for x

OpenStudy (amonoconnor):

X = (27 - 10) = ( 17)

OpenStudy (anonymous):

hmmm ok , so u just subtract ?

OpenStudy (amonoconnor):

Yep! :)

OpenStudy (anonymous):

wait so what if its x-12=7 ? or x-3 1/2=-8

OpenStudy (amonoconnor):

#1 X - 12 = 7 +12 ___________ X = 19 #2 X - 3.5 = (-8) +3.5 ____________ X = -4.5

OpenStudy (anonymous):

1. 19+x=9 2. 1/2+x=-2 3. 16-x=8 4. x+1/3=9 5. -8.7+x=12 6. x-0.8=-2.3 7. 5-x=20 1/4 8. x-16.6=-3.4 9. -9 2/3-x=-1 1/3

OpenStudy (amonoconnor):

Holy pellet.... Okay. Here. Go :)

OpenStudy (anonymous):

lol , sorry . i was never in algebra 2 and my teacher expects me to just know this all

OpenStudy (amonoconnor):

D you need all the work typed out, or do you trust me?

OpenStudy (anonymous):

i trust u , i have 124 pages of math to do , i mean i wont make u help me with all of it less u wanna , but i dont think my teacher wants to check every single problem lol .

OpenStudy (amonoconnor):

1. X = -10 2. X = -2.5 3. X = -8 4. X = 8 2/3 (8.6666) 5. X = 20.7 Here's some :)

OpenStudy (anonymous):

haha awesome ok (: thank u

OpenStudy (amonoconnor):

6. X = -1.5 7. X = 15.25 (15 1/4) 8. X = 13.2 9. X = -8 1/3 (-8.333)

OpenStudy (anonymous):

sweet , ur helping me a lot . ive been stuck on this for like a month

OpenStudy (amonoconnor):

Haha, no problem. ;)

OpenStudy (anonymous):

u wanna still help ?

OpenStudy (amonoconnor):

Bring it ;)

OpenStudy (anonymous):

write how you can use the rule for multiplication or division to solve the equations: 1. 6x=12 multiply by ____ or divide by ____ 2 1/3x=15 multiply by ____ or divide by ______ 3. -7x=21 multiply by ____ or divide by _____ 4. 2/3x=15 multiply by ____ or divide by ____ 5. -1/2x=2 multiply by ____ or divide by ____

OpenStudy (anonymous):

solve for x: 1. 9x=36 2. 10x=50 , which wouldnt x be 5 ? 3. 18x=6 4. -7x=49 5. -8x=-4 6. 1/4x=2 7. -3/2x=12 8. 1/10x=45 9. 1/1,000x=8.35 10. -1/4x=1/2

OpenStudy (amonoconnor):

1. 2, 6 2. 45, 1/3 3. -3, -7 4. 7.5, 2/3 5. -4, -1/2

OpenStudy (anonymous):

ok , next ?

OpenStudy (amonoconnor):

I'm on it ! :) sorry!

OpenStudy (anonymous):

lol its ok

OpenStudy (amonoconnor):

1. 4 2. 5 3. 1/3 4. -7 5. .5 6. 8 7. -8 8. 450 9. 8350 10. -2

OpenStudy (amonoconnor):

:) lols, do I get paid? ;)

OpenStudy (anonymous):

uhmmmm ... tf , i thought u were just doing it outta the kindness of ur heart . guess i thought wrong .

OpenStudy (amonoconnor):

Hey, do you have a webcam on whatever you're using? (Tablet, iPad, PC, laptop... Etc.)

OpenStudy (amonoconnor):

Lols. I am! But if you send me pictures of the problems, I can write on the picture on my iPad! and it will be a lot faster.

OpenStudy (amonoconnor):

amon.oconnor@icloud.com

OpenStudy (anonymous):

no i dont , my camera is broke .

OpenStudy (amonoconnor):

*sadness* Alright, any more :)

OpenStudy (anonymous):

1. 8 times what number equals 24? 2. 12 times what number is equal to 4? 3. 25 times what number equals 5? 4. 1/3 of what number is 9? 5. 2/5 of what number is 1?

OpenStudy (amonoconnor):

1. 3 2. (1/3) or .333 3. (1/5) or .2 4. 27 5. 2.5

OpenStudy (amonoconnor):

I sincerely regret to inform you that I'm going to bed in fifteen minutes :/

OpenStudy (anonymous):

1. 5x+3=18 2. 30-5x=0 3. 17-3x=11 4.22x+5=93 5.7+4x=11 6. 3+6x=21 7. 4x-10=26 8. -3x+7=-14 9. 1/2x-12=16 10. 35-2/3x=13 11. -1/10x+8=46 12. 4/5x+4=8 13. -2/7x-6=-4 14. -9x-9=-9 15. -5/6x-3=27 16. 5/3x-18=-43

OpenStudy (anonymous):

noooo why ?

OpenStudy (amonoconnor):

Alright... I'll stay up ;)

OpenStudy (amonoconnor):

Here I go...

OpenStudy (anonymous):

one step is missing in the solution to each question. using a complete sentance, write the missing step: 1. 8x-18=46 step one: add 18 to both sides of the equation. step two: ________________________________ 2. -1/4x+16=-4 step one: subtract 16 from both sides of the equation. step two: _____________________________________ 3. 1/5x-12=18 step one: _____________________________________ step two: multiply both sides of the equation by 5(or divide by 1/5) 4. 2/3x+6=24 step one: subtract 6 from both sides of the equation step two:_______________________________________

OpenStudy (amonoconnor):

1. 3 2. 6 3. 2 4. 4 5. 1 6. 3 7. 9 8. 7 9. 56 10. 33 11. -380 12. 5 13. -7 14. 0 15. -36 16. -15

OpenStudy (amonoconnor):

1. Divide both sides of the equation by 8. 2. Either divide both sides by the quantity "negative one-quarter", or multiply both sides of the equation by negative four. 3. Add 12 to both sides of the equation. 4. Divide both sides of the equation by the quantity two-thirds.

OpenStudy (amonoconnor):

Looking good to you? Is this helpful?

OpenStudy (anonymous):

ya very

OpenStudy (amonoconnor):

Good :)

OpenStudy (amonoconnor):

Is there more?? ;)

OpenStudy (anonymous):

yupp . lol . hold up lemme try something

OpenStudy (amonoconnor):

Okay

OpenStudy (anonymous):

there

OpenStudy (amonoconnor):

So much nicer. ;)

OpenStudy (amonoconnor):

1. D 2. B 3. E 4. G 5. F 6. C 7. A

OpenStudy (anonymous):

ok next ?

OpenStudy (amonoconnor):

How do you insert a picture to these messages?

OpenStudy (anonymous):

its says attach files

OpenStudy (amonoconnor):

Do you have an email ?

OpenStudy (amonoconnor):

I'm not seeing the link to attach a file :/

OpenStudy (amonoconnor):

I have all the number lines drawn out and everything... :(

OpenStudy (anonymous):

meowimacat96@outlook.com

OpenStudy (amonoconnor):

It'll be coming your way !

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

got it

OpenStudy (amonoconnor):

Sweet. Does that work ? Is that what you needed?

OpenStudy (anonymous):

yeah

OpenStudy (amonoconnor):

I'm not going to be able to stay awake much longer :{

OpenStudy (amonoconnor):

Soo... Tired... XP

OpenStudy (anonymous):

... :(

OpenStudy (anonymous):

i got more tho and have an hour

OpenStudy (amonoconnor):

I'm sorry Savannah2D0pe...

OpenStudy (amonoconnor):

What ?

OpenStudy (anonymous):

i only have an hour to get more math done and i need to get it done , soon . cuz blah , can u do a couple more pages and help me tomorrow night or something ?

OpenStudy (amonoconnor):

*sigh* yeah... Meh sure... ;)

OpenStudy (anonymous):

ok thank you

OpenStudy (amonoconnor):

Lols only for you Savannah... Only for you haha

OpenStudy (anonymous):

sweeett

OpenStudy (anonymous):

write two equations that you need to solve the equations: 1. |x+3|=6 ___________ or ___________ 2. |x-9|=5 __________ or ___________ 3. |x-1/2|=4 _________ or ____________ 4. |x+3 1/4|=7 3/4 _____________or ______________ 5. |x+1.3|=8.5 ____________ or _____________

OpenStudy (amonoconnor):

Okay... 1. E 2. B 3. G 4. D 5. A 6. F 7. C

OpenStudy (anonymous):

wooo !

OpenStudy (amonoconnor):

Haha, how could I resist a cute chick with the screen name savanna ;) lol. Okay: _< Is "less than or equal to" < just "less than" 8. -2 _< x _< 2 9. -3 < x < 1 10. -1 < x < 0 11. -8 _< x _< -5 12. X > 6, x<6 13. X > 2, x < -4 14. X _> -5, X _< -8 15. X > 1, X < -1

OpenStudy (anonymous):

ok sweet

OpenStudy (anonymous):

write two equations that you need to solve the equations: 1. |x+3|=6 ___________ or ___________ 2. |x-9|=5 __________ or ___________ 3. |x-1/2|=4 _________ or ____________ 4. |x+3 1/4|=7 3/4 _____________or ______________ 5. |x+1.3|=8.5 ____________ or _____________

OpenStudy (amonoconnor):

I think this is what they want for these... 1. X = 3 X = -3 2. X = 14 X = -14 3. X = 4.5 X = -4.5 4. X = 4.5 X = -4.5 5. X = 7.2 X = -7.2

OpenStudy (anonymous):

-3+3 doesnt equal 6 tho ?

OpenStudy (amonoconnor):

Oh, I was subtracting from inside the absolute value bracket to the other side. Meh! Idk...

OpenStudy (anonymous):

there saying + or - what would equal 6 when + or - 3

OpenStudy (anonymous):

like 3+3=6 nd 9-3=6

OpenStudy (anonymous):

i just dont know the rest

OpenStudy (amonoconnor):

Yeah, like plus or minus every one, that's the absolute value of x alone.the solution for each one is actually plus or minus one number. Like 3 for the first one.